OpenStudy (anonymous):

Consider (4x + 3y)dx + (2xy)dy = 0 Find n, so that when the differential equation is multiplied by x^n, it becomes exact.

OpenStudy (anonymous):

Math 201 UVic, eh? If so you forgot the y^2, I'll solve assuming it's there. Choose M = 4x + 3y^2 N = 2xy Multiply both by x^n and let My = Nx M = 4x^(n+1) + 3(x^n)y^2 N = 2(x^(n+1))y My = 6(x^n)y = Ny = 2(n+1)(x^n)y 6(x^n)y = 2(n+1)(x^n)y x^n and y cancels on both sides 6 = 2(n+1) 3 = n + 1 n = 2 Just multiply the DE by x^2 and solve like any other exact differential equation.

OpenStudy (anonymous):

You should get (x^3)(y^2) + x^4 = c for an answer.

OpenStudy (anonymous):

Have you figured out 5 and 6 yet?

OpenStudy (anonymous):

haha yeah 201. yeah its not bad once you know the typo. working on 5 and 6 now, have no idea for 5